# Tension-induced giant actuation in unstructured elastic sheets

### v1

@@ Abstract
Buckling is normally associated with a compressive load applied to a slender structure; from railway tracks in extreme heat Pucillo (2016) to microtubules in cytoplasm Brangwynne et al. (2006), axial compression is relieved by out-of-plane buckling. However, recent studies have demonstrated that tension applied to structured thin sheets leads to transverse motion that may be harnessed for novel applications. For instance, tension in one direction can actuate motion in an orthogonal direction, enabling the creation of kirigami grippers Yang et al. (2021), ‘groovy metasheets’ for multi-stable morphing Meeussen and van Hecke (2023), and lateral bending of ribbed sheets Siéfert et al. (2021). Qualitatively similar behaviour has also been observed in simulations of thermalized graphene sheets Chen et al. (2022); Valenzuela et al. (2023), where clamping along one edge leads to tilting in the transverse direction. Here, we show that this counter-intuitive behaviour is, in fact, generic for thin sheets that have a relatively low stretching modulus compared to the bending modulus, which allows ‘giant actuation’ with moderate strain. Indeed, transverse buckling can be induced by uni-axial tension even in unstructured elastic sheets, provided that the axial loading is sufficiently close to localized. We therefore refer to this as ‘Tension indUced Giant actuation’ (TUG actuation) and study its properties. We show that TUG actuation occurs because of an efficient transfer of applied tensile load into compression and determine scaling results for the transverse angle as a function of applied strain; our scaling results compare favorably with both experiments and simulations. Our findings suggest that controlled buckling in tension can be utilized in a broader range of materials and structures to give TUG actuation, potentially expanding the scope of its application in material science and engineering.
@@
Buckling in compression is the archetype of elastic instability Reis (2015): when compressed along its longest dimension, a thin structure such as a playing card will buckle out-of-plane accommodating the imposed compression without a significant change of length. Given its ubiquity, it is unsurprising that buckling is used as the simplest ingredient through which to understand the failure of structures from bridges to marine and aerospace structures Jones (2006). However, recently examples have emerged in which an applied tension can lead to a large out-of-plane deformation — the hallmark of buckling. The first examples of this ‘buckling in tension’ required ingenious design of structures Zaccaria et al. (2011), but have also emerged in natural systems such as the luxation of human fingers Fraldi et al. (2023).
@@
At the same time, there has been increased interest in mechanical meta-materials Bertoldi et al. (2017); Jiao et al. (2023): elastic objects with a structure that allows for novel functionality. Several examples of these meta-materials seem to buckle in an orthogonal direction under uni-axial extension, and are shown in Fig. 1a,b. For example, groovy metasheets (Fig. 1a) Meeussen and van Hecke (2023) and kirigami grippers (Fig. 1b) Yang et al. (2021) look superficially different but both exhibit significant buckling upon a small axial extension. Similar behaviour can be seen in other systems too, including the ribbed sheet Siéfert et al. (2021).
@@
As a first clue that this behaviour may not need the careful design that we might first assume, the reader is invited to crumple a piece of paper, open it out again, clamp the centre between thumb and finger and then pull along one axis. As shown in Fig. 1c, this informal experiment suggests a similar phenomenology: the paper tilts perpendicular to the direction of pulling. A more formal version of this phenomenon has also recently been observed in simulations of (thermal) graphene when one edge is clamped Chen et al. (2022); Valenzuela et al. (2023). This ubiquity of the tensile buckling in these systems suggests something more fundamental that seems not to have been reported previously, and that hence deserves further investigation. A clue on the key ingredients may be understood by modifying the crumpled paper example just considered: if instead of pulling at the centre with fingers, the crumpled paper is firmly clamped in the hands and pulled. Now no tilting is observed.
@@
Figure 1: Uniaxial, localized tension leads to transverse buckling in a range of structured and unstructured thin sheets. a, A groovy plastic sheet is stretched by point force along the horizontal direction and forms a scar along which the sheet bends and folds (left). Diverse morphologies can be realized (right). Adapted from Meeussen and van Hecke (2023). b, A soft gripper made of a kirigami shell is actuated to grasp a raspberry by stretching along the midline. Adapted from Yang et al. (2021). c, A crumpled paper sheet is stretched at the center point along the horizontal direction and demonstrates folding along the transverse direction. d, Image showing the experimental setup and associated geometrical parameters of the sheet. e, Visualizations of the folding angle, $\alpha$, that is observed when the sheet is stretched: left shows the deflection from horizontal of a projected laser line while right shows the deflection of pins attached to the surface which cease being parallel upon deformation. f, Experimentally measured folding angles $\alpha$ as a function of the mean tensile strain, $\bar{\epsilon}$, defined in (1). Plotting on a doubly-logarithmic scale for sheets of various dimensions reveals the power-law relationship $\alpha\propto\bar{\epsilon}^{3/4}$, which we seek to explain in this article. For all specimens shown here, we fix $W_{\mathrm{load}}/W=0.25$.
@@
This observation motivates only using a localized pulling force. Indeed, similar phenomenology can be observed by taking a uniform sheet and stretching with a small clamped region at the centre (Fig. 1d). (In the case of uniform clamping along the edge, the sheet remains planar (see Fig. A2b) suggesting that this behaviour occurs at much smaller strains than are required to cause global wrinkling Cerda et al. (2002); Cerda and Mahadevan (2003).) This phenomenology cannot, therefore, be caused by the structure inherent in the examples already discussed (and shown in Fig. 1a-c). We focus thus on experimental realizations in which a small region along the axis is loaded in sheets that are otherwise uniform.
In our experiments, we use rectangular sheets of polyvinyl chloride (PVC) with uniform thickness $t$ chosen in the range $40\mathrm{~{}\mu m}\leq t\leq 150\mathrm{~{}\mu m}$; the sheets were laser cut with length $L$ and width $W$ and an additional ‘tab’ (of width $W_{\mathrm{load}}=10\mathrm{~{}mm}$ and length $L_{\mathrm{load}}=30\mathrm{~{}mm}$) on each side to ensure that a uniform clamping could be applied. The dimensions of the rectangular sheet (excluding the tab) were varied so that $L\in\{40,80,200\}\mathrm{~{}mm}$ with $W=40\mathrm{~{}mm}$ was fixed, see Fig. 1f. (These dimensions ensure that the resulting sheet has $L/W\geq 1$ with $W/t\gg 1$.) The Young’s modulus of the PVC used here is given by the manufacturer as $E=2-4\mathrm{~{}GPa}$ (measured by tensile test, as shown in Fig. A1c), though its precise value only scales the forces required to cause a given deformation, not the deformation itself.
The rectangular sheets were clamped by the tabs in 3D-printed fixtures that were attached onto mechanical sliders (Fig. 1d and Fig. A1). They were loaded along the $y$-axis by a distance $\Delta u_{y}$. If the sheet remained planar, this imposed displacement would lead to a planar strain $\epsilon_{\mathrm{planar}}=\Delta u_{y}/L$. However, once the strain reached a small threshold value of $\epsilon_{\mathrm{planar}}$ (to be discussed in due course) it curls in both $x$ and $y$ directions. Figure 1e shows this curving as viewed with laser lines and also with small cannula needles attached to the edge of the sheet; these needles allow us to measure the angle, $\alpha$, that the edge of the sheet makes with the horizontal. The evolution of the angle $\alpha$ with additional strain is shown in Fig. 1f; here the value of the strain is measured using a strain gauge placed at the centre of the sample (see Methods for more details). This gives a measure of the local strain $\epsilon_{\mathrm{loc}}$.
@@
A plot of the measured inclination angle, $\alpha$, as a function of the measured strain, $\epsilon_{\mathrm{loc}}$, (see Fig. 1f) shows that $\alpha\propto\bar{\epsilon}^{3/4}$. This scaling is novel in two important respects: firstly, we reiterate that $\epsilon_{\mathrm{loc}}>0$ — the sheet is subject to a tensile load and yet appears to buckle in the transverse direction. Secondly, this scaling is more sensitive to the imposed end-stretching than the regular (compressive) buckling in which, for example, the amplitude $A\propto\Delta^{1/2}$ where $\Delta$ is the imposed end-shortening Howell et al. (2008).
To test whether this behaviour is indeed generic, or is caused by some imperfection or anisotropy of the experimental system, we also performed systematic numerical simulations using the finite element method (FEM) software ABAQUS. The results of our FEM simulations are given in Fig. 2, showing the observed inclination angle as a function of the mean strain
$\bar{\epsilon}=(L^{\prime}-L)/L,$ (1)
where $L^{\prime}$ is the total deformed length. These results demonstrate that the key phenomenology is similar: out-of-plane bending occurs transverse to the imposed stretching (Fig. 2a), while the angle of inclination, $\alpha$, follows a power-law similar to that observed experimentally. One important difference between experiments and these numerical simulations is that we observe a small, and $\bar{\epsilon}$-independent, angle for very small stretching with a very sudden transition to a larger, $\bar{\epsilon}$-dependent, $\alpha$, at some critical strain, $\bar{\epsilon}_{c}$. We therefore seek to understand both the emergence of a critical strain, $\bar{\epsilon}_{c}$, and the scaling of the angle $\alpha$ with $\bar{\epsilon}>\bar{\epsilon}_{c}$.
@@
Figure 2: Finite element simulations of tensile stretching induced folding in elastic sheets. a, Cross-sectional profiles of a sheet stretched along the $y$-axis in normalized units. The color bar indicates the tensile strain of the midline at $x=0$ (see top-left inset). The top-right insets (i-iii) show the deformed shapes of the elastic sheet at increasing strains, with corresponding numbers labelled in the cross-sectional profiles, and colors indicating the $\zeta$-displacements according to the color bar for $U3$. The model shown here has the geometry $L/W=2,t/W=2.5\times 10^{-3},W_{\mathrm{load}}/W=0.1$. b, Simulation results of tensile strains and folding angles for elastic sheets with various geometric parameters.The color bar indicates the thickness-to-width ($t/W$) ratio. The strains and folding angles also display a slope of ${3}/{4}$, illustrated by the dotted guideline.
@@
Our experiments and numerical simulations show a wide range of edge angles $0.01^{\circ}\lesssim\alpha\lesssim 100^{\circ}$ at small imposed strains — the axial strain is at most $\bar{\epsilon}\sim 10^{-2}$ in Figs. 1 and 2. The conjunction of large out-of-plane deflections and small strains is a common feature of thin sheets and is encapsulated in the Föppl-von Kármán (FvK) plate theory, which combines a mechanically linear response (i.e. a Hookean stress-strain relationship) with nonlinear displacement-strain relationships (i.e. out-of-plane displacements contribute to strains).
@@
The FvK equations therefore couple the stress field within the sheet to the sheet’s shape. However, this coupling makes analytical progress difficult. A key simplification comes while the sheet remains planar: we then need only solve for the stress within the loaded sheet. This can be posed as the bi-harmonic equation for the Airy stress function $\chi$ with appropriate boundary conditions (see Appendix D). The key result of this analysis is that, for sufficiently slender sheets at least, and for $W_{\mathrm{load}}\ll W$, the stress components within the sheet remain tensile except within a small region close to the edges (i.e. within a typical distance $W$ from the short edges the transverse stress is compressive). The linearity of this planar problem means that the magnitude of this compressive stress must be linear in the applied tension $T$. The coincidence of lengths (i.e. that the lateral scale of the compressed area scales with $W$) and of stresses (i.e. that the resulting transverse compression is linear in the applied tension, $T$) simplifies the near-threshold analysis of instability Xin and Davidovitch (2021a) so that the classic Euler-like buckling analysis is helpful (though see Appendix D for details). Hence, we expect that buckling will occur when $T$ reaches some multiple of $B/W^{2}$, which immediately gives us that
$\bar{\epsilon}_{c}\propto\frac{B}{YW^{2}},$ (2)
@@
where $Y=E\,t$ is the stretching modulus. For the unstructured, isotropic sheets used here, which have Poisson ratio $\nu$, $B=Yt^{2}/[12(1-\nu^{2})]$ and so (2) predicts that
$\bar{\epsilon}_{c}\propto\frac{t^{2}}{(1-\nu^{2})W^{2}},$ (3)
@@
which may be interpreted as the reciprocal of the von Kármán number Košmrlj and Nelson (2017a). We shall see later that this prediction is borne out by detailed numerical simulations.
Beyond the buckling threshold, we still impose $\Delta u_{y}/L$ but the mean strain $\bar{\epsilon}$ is not known a priori. Since there is additional length in the curved part, we must have that $\bar{\epsilon}>\Delta u_{y}/L$; we observe in our numerics that for $\bar{\epsilon}>\bar{\epsilon}_{c}$ this ‘excess strain’ is proportional to $\bar{\epsilon}-\bar{\epsilon}_{c}$ (see Appendix C). This gives a first physical understanding of the root of buckling: for $\bar{\epsilon}<\bar{\epsilon}_{c}$ the whole planar sheet is subject to significant strain, even though it is not directly loaded (since $W_{\mathrm{load}}/W\ll 1$). Buckling out-of-plane allows a narrow portion of the sheet around the loaded part to accommodate its load, without additionally stretching the unloaded part.
@@
We also observe that the sheet is doubly curved in a narrow region; outside this region, there is a small curvature in the $y$-direction (along the pulling direction), while in the $x$-direction, the sheet is essentially flat (albeit inclined at the angle $\alpha$, see Appendix D). We denote the (unknown) width of this region by $\omega$. To determine an estimate for $\omega$, we use the experimental (and numerical) observation of the shape to pose as an ansatz for the pulled sheet’s shape
@@
$\zeta(x,y)=\delta\times\begin{cases}\sin\frac{\pi y}{L}\cos\frac{\pi x}{\omega},\quad|x|<\omega/2,\\ \frac{\pi}{\omega}\left(\frac{\omega}{2}-|x|\right),\quad\omega/2\leq|x|\leq W/2,\end{cases}$ (4)
@@
where $\delta$ is the maximum amplitude, and the constants have been chosen to ensure that the vertical displacement, slope and curvature of the sheet are all continuous at the edge of the curved central region. Note that this ansatz is only appropriate for $W_{\mathrm{load}}\ll\omega$.
@@
The width of the curved region $\omega$ is not yet known. To determine this, we note that the smaller the value of $\omega$ the more curved the sheet is in the central region, and the higher its bending energy is: $U_{\mathrm{bend}}\sim B\int~{}(\nabla^{2}\zeta)^{2}\mathrm{d}A\sim B\delta^{2}L/\omega^{3}$ (where we have neglected the bending energy due to the curvature in the $y$-direction because $L\gtrsim W\gg\omega$). Conversely, the excess stretching energy caused by being doubly-curved is localized to this central region so that $U_{\mathrm{stretch}}\sim\int\sigma:\epsilon~{}\mathrm{d}A\sim WLT^{2}/Y+\omega LT\delta^{2}/L^{2}$, after we make the common assumption that the sheet buckles in the compressing ($x$) direction to relax the compressive stress. Notice also that there are two different length scales for the two terms in $U_{\mathrm{stretch}}$: the whole sheet, of width $W$, feels the longitudinal tension, but only the narrow strip, of width $\omega$, is subject to the additional strain, $\sim\delta^{2}/L^{2}$, from double-curvature. In the first term in $U_{\mathrm{stretch}}$ we also assume that $\Delta u/L\sim T/Y$ remains true after buckling, by virtue of the observation that $\bar{\epsilon}-\Delta u_{y}/L\propto\bar{\epsilon}-\bar{\epsilon}_{c}$ sufficiently close to the buckling threshold ($(\bar{\epsilon}-\bar{\epsilon}_{c})/\bar{\epsilon}_{c}\lesssim 10$, from Fig. A5b). The total elastic energy, $U_{\mathrm{bend}}+U_{\mathrm{stretch}}$, is then minimized when
$\omega\propto\left(\frac{BL^{2}}{T}\right)^{1/4}=\left(\frac{BL^{2}}{Y\bar{\epsilon}}\right)^{1/4}.$ (5)

### v2

# Localized tension-induced giant folding in unstructured elastic sheets

@@ Abstract
Buckling in compression is the archetype of elastic instability: when compressed along its longest dimension, a thin structure such as a playing card will buckle out-of-plane accommodating the imposed compression without a significant change of length. However, recent studies have demonstrated that tension applied to sheets with microscopic structure leads to out-of-plane deformation in applications from ‘groovy metasheets’ for multi-stable morphing to kirigami grippers. Here, we demonstrate that this counter-intuitive behavior — a large transverse folding induced by a relatively small imposed longitudinal tension — occurs also in unstructured sheets of isotropic material. The key to this behavior is that a localized uniaxial tension induces giant folding; we refer to this as ‘localized TUG folding’ to reflect the importance of localized tension and its mode of actuation. We show that localized TUG folding occurs because of an efficient transfer of applied tensile load into compression — a geometric consequence of a localized applied tension. We determine scaling results for the folding angle as a function of applied strain in agreement with both experiments and simulations. The generic nature of localized TUG folding suggests that it might be utilized in a broader range of materials and structures than previously realized.
Buckling |||| Metamaterials |||| Linear elasticity
@@
Buckling is associated with compressive loads on a slender structure Reis (2015); from railway tracks in extreme heat Pucillo (2016) to microtubules in cytoplasm Brangwynne et al. (2006), axial compression is relieved by out-of-plane buckling. Given its ubiquity, it is unsurprising that buckling under compression is considered a fundamental mechanism through which structures from bridges to marine and aerospace structures may fail Jones (2006); Reis (2015). However, other examples have emerged in which an applied tension can lead to a large out-of-plane deformation — the hallmark of buckling. The first examples of this ‘buckling in tension’ required ingenious design of structures Zaccaria et al. (2011), but are now also understood to exist in natural systems such as the luxation of human fingers Fraldi et al. (2023).
@@
At the same time, there has been increased interest in mechanical metamaterials Bertoldi et al. (2017); Jiao et al. (2023): elastic objects with a structure that allows for novel functionality. Several examples of these metamaterials seem to buckle in an orthogonal direction under uniaxial extension, see Fig. 1a,b. For example, groovy metasheets (Fig. 1a) Meeussen and van Hecke (2023) and kirigami grippers (Fig. 1b) Yang et al. (2021) look superficially different but both exhibit significant transverse motion upon a small axial extension. Similar behavior can be seen in other systems too, including the ribbed sheet Siéfert et al. (2021), while the leaves of some trees fold as senescence progresses (Fig. 1c).
@@
As a first clue that this behavior may not need the careful design that might first be assumed, the reader is invited to crumple a piece of paper, open it out again, clamp the center between thumb and finger and then pull along one axis. As shown in Fig. 1d, this informal experiment suggests a similar phenomenology: the paper folds perpendicular to the direction of pulling. A more formal version of this phenomenon has also recently been observed in simulations of (thermal) graphene in which one edge is clamped, leading to extension along this edge Chen et al. (2022); Valenzuela et al. (2023). This ubiquity of the tensile buckling in these systems (both of which are often modelled as isotropic solids, albeit with modified effective moduli) suggests something more fundamental that has not been reported previously. A clue to the key ingredients may be understood by modifying the crumpled paper example just considered: if instead of pulling at the center with two fingers, the crumpled paper is firmly clamped by hand along its whole width and pulled, no folding is observed.
Perhaps the most striking feature of the examples shown in Fig. 1a-d is that a very small and localized tensile deformation in-plane induces a very large folding angle; for this reason we refer to this means of inducing a fold angle as localized Tension-indUced Giant folding, or localized TUG folding. In this paper, we use a combination of experiments, numerical simulations and mathematical modelling to understand the origin of localized TUG folding.
Through the application of the Föppl-von Kármán (FvK) plate theory Landau and Lifshitz (1986), we examine both the onset of out-of-plane deformation (the near-threshold regime) and its evolution far beyond the critical imposed strain (the far-from-threshold regime Cerda et al. (2002); Cerda and Mahadevan (2003); Davidovitch et al. (2011)). We derive the scaling law for the inclination angle as a function of the applied strain that explains the results in experiments and simulations. This scaling law provides useful insight into the geometry and material properties that allow for a large actuation of an elastic sheet.
@@
Figure 1: Uniaxial, localized tension leads to transverse buckling in a range of structured and unstructured thin sheets. a A groovy plastic sheet is stretched by point force along the horizontal direction and forms a scar along which the sheet bends and folds. Adapted from Meeussen and van Hecke (2023). b A soft gripper made of a kirigami shell is actuated by stretching along the midline to grasp a raspberry. Adapted from Yang et al. (2021). c Leaves of the flowering cherry (Prunus ‘kanzan’) show a similar folding at different stages of leaf senescence from early (top) to mid (bottom). d A crumpled paper sheet is stretched along the horizontal direction by a localized force, and demonstrates folding in the transverse direction. e Schematic of the experimental setup. A rectangular planar sheet is clamped at the additional ‘tabs’ centered on each short side (upper diagram). The sheet folds transversally, developing a non-zero fold angle $\alpha$, as the clamps are displaced to impose a longitudinal displacement (lower diagram). f Our experimental setup and associated geometrical parameters of the sheet. Side (upper snapshots) and zenithal view (lower snapshots) of the planar, unloaded, sheet (left), and the loaded sheet when it folds in the transverse direction (right). g Experimentally measured inclination angles $\alpha$ as a function of the local tensile strain, $\epsilon_{\mathrm{loc}}$, measured by a strain gauge for a PVC sheet. Plotting on a doubly-logarithmic scale for sheets of various dimensions reveals the power-law relationship $\alpha\propto\epsilon_{\mathrm{loc}}^{3/4}$, which we seek to explain in this article. For all specimens shown here, $W=40\mathrm{~{}mm}$, $W_{\mathrm{load}}/W=1/4$, and other geometric parameters are listed in the legend. The Young’s modulus is approximately $3\mathrm{~{}GPa}$ from tensile test measurements on fully clamped strips (see Fig. S1c in the SI Appendix); the Poisson’s ratio $\nu\approx 0.38$ for PVC materials.
Importance of localized tension
@@
The difference between pulling at a point versus along the full edge motivates only using a localized pulling force. Indeed, similar phenomenology can be observed by taking a uniform sheet and stretching with a small clamped region at the centre (Fig. 1e). (In the case of uniform clamping along the whole edge, the sheet remains planar — see the numerical simulations of a stretched sheet with fully clamped edges in the SI Appendix (Fig. S3b) — suggesting that this behavior occurs at much smaller strains than are required to cause global wrinkling Cerda et al. (2002); Cerda and Mahadevan (2003).) This phenomenology cannot, therefore, be caused only by the structure, cuts or folds, in the examples already discussed (Fig. 1a-d). We therefore focus on experimental realizations in which a small region along the axis is loaded in sheets that are unstructured (i.e. spatially homogeneous and isotropic).
In our experiments, we use rectangular sheets of polyvinyl chloride (PVC), with uniform thickness $t$, that are laser cut with length $L$ and width $W$ and an additional ‘tab’ (of width $W_{\mathrm{load}}=10\mathrm{~{}mm}$ and length $L_{\mathrm{load}}=30\mathrm{~{}mm}$) on each side to ensure that a localized uniform clamping could be applied. The dimensions of the rectangular sheet (excluding the tab) are varied ensuring that the resulting sheet has $L/W\geq 1$ with $W/t\gg 1$, see Fig. 1f. (Further details of the sheets used in experiments may be found in Methods and the SI Appendix; images of our experimental setup are shown in Fig. S1.) The rectangular sheets are clamped by the tabs in 3D-printed fixtures that are attached onto mechanical sliders (Fig. 1f). The supports are displaced along the $y$-axis by a distance
$\Delta u_{y}\equiv u_{y}(x=0,y=L/2)-u_{y}(x=0,y=-L/2)>0.$ (1)
If the sheet remained planar, so that its mid-plane $z=0$ (see Fig. 1f), this imposed displacement would lead to a planar strain
$\epsilon_{0}=\frac{\Delta u_{y}}{L}.$ (2)
However, once the strain reaches a small threshold value of it curls in both $x$ and $y$ directions so that the mid-plane of the sheet is deformed to $z=\zeta(x,y)$. Double curving is displayed by a projected laser line along the two midlines at $x=0$ and at $y=0$ — images are included in Fig. S1e,f of the SI Appendix — and distinguishes this effect from the edge clamping-induced tilting of graphene Chen et al. (2022); Valenzuela et al. (2023), in which the imposed boundary condition inhibits this double curving. Figures 1e,f show the transverse curving as viewed with small cannula needles attached to the edge of the sheet; these needles allow us to measure the angle, $\alpha$, that the edge of the sheet makes with the horizontal.
Scaling behavior of folding
@@
The angle, $\alpha$, that the edge of the sheet makes with the horizontal is measured as a function of imposed strain, see Fig. 1g. (Here the value of the strain is measured using a strain gauge placed at the center of the sample, which gives a measure of the local strain $\epsilon_{\mathrm{loc}}$ — see Methods.) The experimental results of Fig. 1g suggests that $\alpha\propto\epsilon_{\mathrm{loc}}^{3/4}$ — a scaling that is also consistent with similar experiments on crumpled paper (see Fig. S2 of the SI). This scaling is novel in two important respects: firstly, we reiterate that $\epsilon_{\mathrm{loc}}>0$ — the sheet is subject to a tensile load and yet appears to buckle in the transverse direction. Secondly, this scaling is more sensitive to the imposed end-stretching than the regular (compressive) buckling in which, for example, the amplitude $A\propto\Delta u_{y}^{1/2}$ Howell et al. (2008) where $\Delta u_{y}$ is the imposed end-shortening defined in (1).
@@
Figure 2: Finite element simulations of tensile stretching induced folding in elastic sheets. a Cross-sectional profiles, $\zeta(x,0)$, of a sheet stretched along the $y$-axis in normalized units. The color bar indicates the tensile strain of the midline at $x=0$ (see top-left inset). The top-right insets (i-iii) show the deformed shapes of the elastic sheet at increasing strains, with corresponding numbers labelled in the cross-sectional profiles, and colors indicating the $\zeta$-displacements normalized by $W$ according to the color bar. The model shown here has the following geometric and material parameters: $L/W=2,t/W=2.5\times 10^{-3},W_{\mathrm{load}}/W=1/10$, $E=3\mathrm{~{}GPa}$, and $\nu=0.3$. b Simulation results of tensile strains and inclination angles for elastic sheets with various geometric parameters and Poisson’s ratio, as indicated in the legend. The color bar indicates the thickness-to-width ($t/W$) ratio. These numerical results suggest that $\alpha\propto\bar{\epsilon}^{3/4}$ (dashed line), as also observed experimentally. Note that in (b) the marker style is used to encode the geometry and Poisson’s ratio used in simulations (as indicated by the grayscale legends). Color is used to encode the thickness (as per the colorbar on the right hand side).
To test whether this behavior is indeed generic, or is caused by some imperfection or anisotropy of the experimental system, we also performed systematic numerical simulations using the finite element method (FEM) software ABAQUS. The results of our FEM simulations allow the deformed shape $\zeta(x,y)$ to be calculated. In particular, the observed inclination angle is plotted as a function of the mean strain
$\bar{\epsilon}=\epsilon_{0}+\frac{1}{2L}\int_{-L/2}^{L/2}\left(\frac{\partial\zeta}{\partial y}\right)^{2}~{}\mathrm{d}y,$ (3)
in Fig. 2b. These results demonstrate that the key phenomenology is similar: out-of-plane bending occurs transverse to the imposed stretching (see Fig. 2a, which displays $\zeta(x,0)$, the vertical displacement along the midline at $y=0$), while the angle of inclination, $\alpha$, follows a power-law similar to that observed experimentally, cf. Fig. 1g. Numerical simulations allow us to test the role of the Poisson’s ratio $\nu$; our results show a small effect of $\nu$ but, crucially, that the effect persists as $\nu\to 0$ (see Fig. 2b). This is in contrast with the tensional wrinkling experiment of Cerda and Mahadevan (CM) Cerda and Mahadevan (2003) for which the effect disappears entirely as $\nu\to 0$ Xin and Davidovitch (2021a, b). We further discuss the role of Poisson’s effect in the subsequent theoretical analysis.
One important difference between experiments and these numerical simulations is that we observe a small (experimentally-undetectable), and $\bar{\epsilon}$-independent, angle for very small stretching with a sharp transition to a larger, $\bar{\epsilon}$-dependent, $\alpha$, at some critical strain, $\bar{\epsilon}_{c}$. Note that this transition is not observable in experiments because of the limited angular resolution, which is caused by small fluctuations in the initial state. We therefore seek to understand both the emergence of a critical strain, $\bar{\epsilon}_{c}$, and the scaling of the angle $\alpha$ with $\bar{\epsilon}>\bar{\epsilon}_{c}$.
@@
Our experiments and numerical simulations show a wide range of inclination angles $0.01^{\circ}\lesssim\alpha\lesssim 100^{\circ}$ at small imposed strains — the axial strain is at most $\bar{\epsilon}\sim 10^{-2}$ in Figs. 1 and 2. The conjunction of large out-of-plane deflections and small strains is a common feature of thin sheets and is encapsulated in the FvK plate theory, which combines a mechanically linear response (i.e. a Hookean stress-strain relationship) with nonlinear displacement-strain relationships (i.e. out-of-plane displacements contribute to strains) Landau and Lifshitz (1986). The coupling of the stress field within the sheet to the sheet’s shape in the FvK equations makes analytical progress difficult. A key simplification comes while the sheet remains planar: we then need only solve for the stress within the loaded sheet and the characteristic tension is linear in the imposed end separation,
$T_{0}=Y\epsilon_{0},$ (4)
where $Y=E\,t$ is the stretching modulus; $E$ denotes the Young’s modulus, which is measured experimentally (see SI), but does not directly control this geometrical effect. The problem for the planar stress can be posed as the bi-harmonic equation for the Airy stress function $\chi$ with appropriate boundary conditions (see the mathematical model in the SI Appendix). The resulting equations can be analyzed using Laplace transform methods Williams (1952); Benthem (1963); Morse and Feshbach (1981) (see SI Appendix for details). The key result of this analysis is that, for sufficiently slender sheets at least, and for $W_{\mathrm{load}}\ll W$, the stress components within the sheet remain tensile except within a small region close to the edges (i.e. within a typical distance $W$ from the short edges the transverse stress is compressive) (see Fig. S9). Moreover, this analysis allows us to compare the effect of point-loading, $W_{\mathrm{load}}\ll W$, with the more common full-width loading, $W_{\mathrm{load}}=W$, in the limit of vanishing Poisson ratio $\nu\to 0$. This comparison (see Fig. S12 of the SI Appendix) shows that the zone of compression persists with point-loading but disappears with full-width loading — the geometry of the loading is key to the localized TUG-folding demonstrated here and distinguishes it from the tensional wrinkling studied extensively previously Cerda and Mahadevan (2003). In addition, this effect does not require a negative Poisson ratio, as might be inferred on the basis of the tilt phase in clamped thermal graphene Chen et al. (2022). The origin of this compression signals the importance of a ‘mechanism’ in generating the compression (see Fig. S12 of the SI Appendix).
@@
The linearity of this planar problem for the stress prior to buckling, means that the magnitude of the compressive stress generated by ‘tugging’ must be linear in the characteristic tension $T_{0}$. The coincidence of lengths (i.e. that the lateral scale of the compressed area scales with $W$) and of stresses (i.e. that the resulting transverse compression is linear in the characteristic tension, $T_{0}$) simplifies the near-threshold analysis of instability Xin and Davidovitch (2021a) so that the classic Euler-like buckling analysis is helpful (see the mathematical model in the SI Appendix for details). Hence, we expect that buckling will occur when $T_{0}$ reaches some multiple of $B/W^{2}$, which immediately gives us that
$\epsilon_{0}^{c}=\bar{\epsilon}_{c}\propto\frac{B}{YW^{2}}.$ (5)
@@
For the unstructured, isotropic sheets used here, the bending modulus $B=Yt^{2}/[12(1-\nu^{2})]$ and so (5) predicts that
$\epsilon_{0}^{c}=\bar{\epsilon}_{c}\propto\frac{t^{2}}{(1-\nu^{2})W^{2}},$ (6)
@@
which may be interpreted as the reciprocal of the von Kármán number Košmrlj and Nelson (2017a). We emphasize that the scalings for the critical strain in (5) and (6) are the standard scalings for Euler buckling Lidmar et al. (2003); Chen et al. (2022) — we shall see later that they are borne out by detailed numerical simulations, which serve to show that this transition is indeed driven by buckling.
@@
We also observe that the sheet is doubly curved in a narrow region (see Fig. 2a); outside this region, there is a small curvature in the $y$-direction (along the pulling direction), while in the $x$-direction, the sheet is essentially flat (albeit inclined at the angle $\alpha$) — in this sense the response of the sheet to uniaxial tension is to fold in the transverse direction. We denote the (unknown) width of this region by $\omega$ (see Fig. 2a) and estimate it using the experimental (and numerical) observation of the shape to pose an ansatz for the sheet’s out-of-plane displacement, $\zeta(x,y)$, relative to the initially planar state; in particular, we take:
@@
$\zeta(x,y)=\delta\times\begin{cases}\cos\frac{\pi y}{L}\cos\frac{\pi x}{\omega},\quad|x|<\omega/2,\\ \frac{\pi}{\omega}\left(\frac{\omega}{2}-|x|\right),\quad\omega/2\leq|x|\leq W/2,\end{cases}$ (7)
@@
where $\delta$ is the maximum amplitude (see Fig. 2a), and the constants have been chosen to ensure that the vertical displacement, slope and curvature of the sheet are all continuous at the edge of the curved central region. (Note that this ansatz is only appropriate for $W_{\mathrm{load}}\ll\omega$, since otherwise a third region would be required.) To determine $\omega$, we note that the smaller its value, the more curved the sheet is in the central region, and the higher its bending energy is: $U_{\mathrm{bend}}\sim B\int~{}(\nabla^{2}\zeta)^{2}\mathrm{d}A\sim B\delta^{2}L/\omega^{3}$ (where we have neglected the bending energy caused by the curvature in the $y$-direction because $L\gtrsim W\gg\omega$). Conversely, the excess stretching energy caused by being doubly-curved is localized to this central region so that $U_{\mathrm{stretch}}\sim\int\sigma:\epsilon~{}\mathrm{d}A\sim WL\bar{T}^{2}/Y+\omega L\bar{T}\delta^{2}/L^{2}$, where the characteristic tension, $\bar{T}$, is linear in the mean strain,
$\bar{T}=Y\bar{\epsilon},$ (8)
@@
and we make the common assumption that the sheet buckles in the compressing ($x$) direction to relax the compressive stress. Notice also that there are two different length scales for the two terms in $U_{\mathrm{stretch}}$: the whole sheet, of width $W$, feels the longitudinal tension, but only the narrow strip, of width $\omega$, is subject to the additional strain, $\sim\delta^{2}/L^{2}$, from double-curvature. In the first term in $U_{\mathrm{stretch}}$ we also assume that $\epsilon_{0}\sim\bar{\epsilon}=\bar{T}/Y$ remains true after buckling, by virtue of the observation that $\bar{\epsilon}-\epsilon_{0}\propto\epsilon_{0}-\bar{\epsilon}_{c}$ sufficiently close to the buckling threshold ($(\bar{\epsilon}-\bar{\epsilon}_{c})/\bar{\epsilon}_{c}\lesssim 10$ — see the numerical simulation results for the mean strain beyond buckling in Fig. S6b in the SI Appendix). The total elastic energy, $U_{\mathrm{bend}}+U_{\mathrm{stretch}}$, is then minimized when
$\omega\propto\left(\frac{BL^{2}}{\bar{T}}\right)^{1/4}=\left(\frac{BL^{2}}{Y\bar{\epsilon}}\right)^{1/4}.$ (9)